N2(g) + 3H2(g)=2NH3(g)

what is the number of molecules of ammonia, NH3 produced if 6.072dm3 of nitrogen gas in STP is reacted completely with hydrogen gas
1
Expert's answer
2021-03-08T05:57:20-0500

V (N2) = 6.072 dm3;

n (N2) = V/Vm; n = 6.072 dm3/22.4 dm3 = 0.27 mole;

if n (N2) = 0.27 mole then n (NH3) = 0.27 * 2 = 1.35 mole;

N (NH3) = NA * n = 6.02*1023 * 1.35 mole = 8.127*1023

Answer: N (NH3) = 8.127*1023